# Multi-turn Jailbreaking via Global Refinement and Active Fabrication

### v1

@@ Abstract
Large Language Models (LLMs) have achieved exceptional performance across a wide range of tasks. However, they still pose significant safety risks due to the potential misuse for malicious purposes. Jailbreaks, which aim to elicit models to generate harmful content, play a critical role in identifying the underlying security threats. Recent jailbreaking primarily focuses on single-turn scenarios, while the more complicated multi-turn scenarios remain underexplored. Moreover, existing multi-turn jailbreaking techniques struggle to adapt to the evolving dynamics of dialogue as the interaction progresses. To address this limitation, we propose a novel multi-turn jailbreaking method that refines the jailbreaking path globally at each interaction. We also actively fabricate model responses to suppress safety-related warnings, thereby increasing the likelihood of eliciting harmful outputs in subsequent questions. Experimental results demonstrate the superior performance of our method compared with existing single-turn and multi-turn jailbreaking techniques across six state-of-the-art LLMs. Our code is publicly available at https://github.com/Ytang520/Multi-Turn_jailbreaking_Global-Refinment_and_Active-Fabrication.
Multi-turn Jailbreaking via Global Refinement and Active Fabrication
[WARNING] This paper contains unsafe model responses. Reader discretion is advised.
Hua Tang1  Lingyong Yan2thanks:  Corresponding author.  Yukun Zhao2  Shuaiqiang Wang2 Jizhou Huang2Dawei Yin2 1Shanghai Jiao Tong University, Shanghai, China 2Baidu Inc., Beijing, China {htang2023126578,lingyongy}@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated superior performance across various tasks Kim et al. (2024); Cheong et al. (2024); Tang et al. (2023). However, they still face the risk of being misused through malicious instructions  Gupta et al. (2023); Chao et al. (2023); Yi et al. (2024a). Therefore, recent studies explore various safety-enhancing mechanisms to help the model avoid following the malicious instructions Xu et al. (2024); Inan et al. (2023); Zeng et al. (2024a). Among these studies, jailbreak Liu et al. (2023); Chao et al. (2023); Li et al. (2024b) typically stands as a red team technique, which aims at bypassing safeguards and exposing safety weaknesses in LLMs, which can then enhance model security based on the discovered vulnerabilities.
Figure 1: An example of multi-turn jailbreaking with ChatGPT, compared with directly requesting the malicious query.
Recent jailbreaking methods on LLMs can be categorized into two types: (1) single-turn jailbreaking Liu et al. (2023); Chao et al. (2023); Li et al. (2024b); Yi et al. (2024a), where the target model is required to be jailbroken using only one prompt. For instance, Zou et al. (2023) optimizes suffix texts to increase the likelihood of the model generating harmful responses, and Ding et al. (2023) rephrases prompts and embeds them within common task scenarios to make them less detectable. However, these approaches typically involve explicit malicious content, which can be easily discovered and defended by the target model Russinovich et al. (2024). Other methods attempt to obscure the malicious content by rephrasing or incorporating complex logical chains, which may mislead the LLMs into generating irrelevant responses Souly et al. (2024); Huang et al. (2024). (2) Multi-turn jailbreaking Russinovich et al. (2024); Li et al. (2024a); Russinovich et al. (2024); Gibbs et al. (2024); Ren et al. (2024) introduces an attack model to jailbreak the target model via multi-turn dialogues, as shown in Figure 1. The attack model is designed to construct a series of jailbreaking queries $Q=\{q_{1},q_{2},...,q_{N}\}$(also denoted as the jailbreaking path), which can conceal the malicious intent but gradually lead to a successful jailbreak. However, previous multi-turn jailbreaking methods mostly utilize pre-defined templates to generate the jailbreaking path Yang et al. (2024a); Jiang et al. (2024), and keep the queries fixed Cheng et al. (2024) throughout the attack. Other methods refine the queries during the attack process, though they typically perform only local updates by modifying the next query. This effort can lead to off-topic queries and may not be sufficient to succeed within the maximum number of attempts Ren et al. (2024); Russinovich et al. (2024). Besides, previous methods simply accept and concatenate the target model’s intermediate responses to the dialogue history without any modification, which may hinder the model’s ability to answer subsequent queries and result in suboptimal performance.
@@
To this end, we propose a novel multi-turn jailbreaking method, which globally refines the jailbreaking path and actively fabricates the dialogue history throughout the multi-turn jailbreaking process. We begin by requesting our attack model to generate an initial comprehensive jailbreaking path based on the given task. At each turn $i$, the query $q_{i}$ in the jailbreaking path $Q$ is first fed into the target model to obtain the corresponding answer $a_{i}$. If the target model refuses to answer the query $q_{i}$, the attack model first revises $q_{i}$ until a successful response is obtained or the maximum number of retries is reached. Otherwise, all subsequent queries $Q_{\geq i}$ are revised. Once a non-rejective answer is returned, the model globally refines all following queries $Q_{>i}$ based on the dialogue history up to $a_{i}$. In contrast to prior work, our approach refines all remaining queries at each step, allowing the model to stay focused on ultimate jailbreak success and to avoid generating off-topic or ineffective queries. We further introduce an active fabrication mechanism with two key strategies: (1) if the target model continues to reject $q_{i}$, we discard $(q_{i},a_{i})$ and proceed to $q_{i+1}$; (2) When a non-rejective answer $a_{i}$ is returned, we proactively modify it by removing any safety-related warning. These strategies enhance the attacking model’s ability to recover from rejections and maintain effective progress toward a successful jailbreak. We repeat the above process until the final query $q_{N}$ receives a valid response.
@@
  • We introduce a novel multi-turn jailbreaking method that globally refines the jailbreaking path through updating all the remaining queries inside the jailbreaking path to adapt to the context, and actively fabricates the dialogue history to enhance the jailbreaking success.
  • We perform extensive analyses of multi-turn jailbreaking methods. We find that our approach shifts the representations of harmful queries closer to those of harmless ones, which helps explain the effectiveness of multi-turn attacks.
@@ 2.1 Responsible Large Language Models
Large Language Models have demonstrated exceptional capabilities across various domains Wei et al. (2022a, b); Kim et al. (2024); Cheong et al. (2024); Tang et al. (2023). However, they can generate responses that violate human values due to the presence of inappropriate data in the unfiltered training corpus Cao et al. (2023); Zou et al. (2023); Yi et al. (2024a). Efforts have been dedicated to aligning LLMs with human values Inan et al. (2023); Rafailov et al. (2023); Korbak et al. (2023); Ji et al. (2023). During these efforts, jailbreaking, whose aim is to test the safety mechanisms within LLMs, is crucial for identifying the underlying safety issues and thus helps to improve their safety capabilities Zou et al. (2023); Yi et al. (2024a); Chao et al. (2023).
@@ 2.2 Jailbreaking
Mainstream attacks primarily focus on single-turn jailbreaks, which allow only a single attack on the target model during the test phase Zou et al. (2023); Yao et al. (2024); Yi et al. (2024a), like using optimization-based methods to generate adversarial prompts Zou et al. (2023); Liu et al. (2023) or manipulate the prompt itself to achieve jailbreaks Ding et al. (2023); Zeng et al. (2024b); Chao et al. (2023) Most of these methods include explicit malicious content to convey their intent to the target model, and can be easily defended by those input filters Russinovich et al. (2024). Some methods attempt to obscure these phrases by rephrasing the harmful queries Huang et al. (2024); Souly et al. (2024), these attempts can confuse the LLM, leading it to generate general or even irrelevant responses that fail to fulfill the user’s intent.
@@
Given the real-world scenario of multiple interactions with LLMs, some researchers have extended attacks to multi-turn jailbreaking. Multi-turn jailbreaking allows multiple attacks on the target model, with its final response during the dialogue evaluated for the success or failure of the attack Russinovich et al. (2024); Gibbs et al. (2024); Li et al. (2024a); Ren et al. (2024); Yang et al. (2024a); Jiang et al. (2024). Therefore, malicious intent can be concealed within the jailbreaking path Russinovich et al. (2024); Ren et al. (2024), rather than being explicitly stated in the jailbreaking queries. For instance,  Russinovich et al. (2024) designs query chains to avoid explicitly expressing harmful content in the final query, while  Wang et al. (2024); Sun et al. (2024) iteratively refines queries to gradually align query semantics with user intent. However, these approaches often rely on fixed templates  Yang et al. (2024a); Jiang et al. (2024). Some methods Ren et al. (2024); Russinovich et al. (2024) attempt to refine queries during the attack, but they typically make only local updates by modifying the immediate next query, which can result in off-topic queries and may fail to jailbreak within the maximum number of tries. In addition, previous approaches often directly append the target model’s responses to the dialogue history without any revision, which may hinder the model from answering subsequent queries and lead to suboptimal performance.
@@
Our approach differs from prior work in two main ways. First, we globally refine the jailbreaking paths through updating all the subsequent queries at once to adapt to the dynamic dialogue history. In contrast,  Ren et al. (2024) only refines the query when it is rejected by the target model, and  Russinovich et al. (2024) generates the next query based solely on the previous turn. Second, instead of simply appending the target model’s responses to the dialogue history, we detect and remove safety-related warnings, which helps the attack model to smoothly progress toward jailbreaking success.
3 Method
@@
In this section, we introduce our proposed method. We begin with the initialization of the jailbreaking path in subsection 3.1. We then present our Global Path Refinement in subsection 3.2, and Active Fabrication techniques in subsection 3.3. The overall workflow is illustrated in Figure 2.
@@
Figure 2: The workflow of our method. The attack model first initialized a jailbreaking path based on the given task. At each turn, the query $q_{i}$ in the jailbreaking path $Q$ is first fed into the target model to obtain the corresponding answer $a_{i}$. If the query $q_{i}$ is refused, the model first revises $q_{i}$ until a successful response is obtained or the retry limit is reached. Otherwise, we revise all subsequent queries $Q_{\geq i}$; Once a non-rejective answer is returned, the model globally refines all following queries $Q_{>i}$ based on the dialogue history up to $a_{i}$ (Denoted by the Red Line). If the target model continues to reject $q_{i}$, the $(q_{i},a_{i})$ will be discarded and the attack model proceeds to $q_{i+1}$; Besides, when a non-rejective answer $a_{i}$ is returned, we modify it by removing any safety-related warning (Denoted by the Blue Line). The above process is repeated until the final query $q_{N}$ receives a valid response.
3.1 Path Initialization
Since LLMs may refuse to generate the jailbreaking path, following Ren et al. (2024), for the jailbreaking goal (e.g, the bomb construction instruction), we begin by querying the attack model for related information (e.g., a well-known criminal Ramzi Yousef who builds bombs for terrorism). Based on the information and the jailbreaking goal, the attack model is requested to generate an initial innocuous query (e.g., What have Ramzi Yousef done in history?), followed by a sequence of queries, which are prompted to guide target model’ responses to become increasingly relevant to the harmful target until reaching it. Further implementation details can be found in Appendix A.1, and the example initialized paths can be referred to Appendix G.6.1.
@@ 3.2 Global Refinement
As shown in Figure 2, during each turn $i$, the query $q_{i}$ in the jailbreaking path $Q=\{q_{1},q_{2},...,q_{n}\}$ is first fed into the target model to obtain the corresponding answer $a_{i}$. Despite the efforts to conceal malicious intent within the jailbreaking path, some queries may still contain harmful content and be rejected by the target model. If a query is rejected, we first revise $q_{i}$ to replace the harmful content with referential expressions such as ’that in your previous response’. This helps to conceal explicit malicious content within the conversation history while preserving the core intent of each query. However, if the query continues to be rejected after the maximum number of attempts, we revise all subsequent queries $Q_{\geq i}$ to redirect the current jailbreaking path. This changes the core intent of $q_{i}$ and further conceals its malicious intent. For simplicity, we use the attack model to determine whether the response answers or rejects the intermediate query during the attack process.
Once the non-rejection answer is obtained, the model then globally updates the remaining queries $Q_{>i}=\{q_{i+1},...,q_{N}\}$ based on the dialogue history up to $a_{i}$. The refinement helps the attack model to discover more effective and stealthy jailbreaking paths with the comprehensive dialogue history. This global refinement strategy, which updates all subsequent queries at each step, helps the model remain focused on achieving a successful jailbreak. Unlike previous approaches Russinovich et al. (2024); Yang et al. (2024b), our method reduces the likelihood of generating off-topic or ineffective queries, increasing the chances that the path will reach the target within the allowed number of interactions. To support this process, we also monitor the final query to determine whether its topic deviates from the target goal. If it happens, we iteratively refine the path until it becomes relevant or the maximum number of attempts is reached.

### v2

# GRAF: Multi-turn Jailbreaking via Global Refinement and Active Fabrication

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable performance across diverse tasks. Nevertheless, they still pose notable safety risks due to potential misuse for malicious purposes. Jailbreaking, which seeks to induce models to generate harmful content through single-turn or multi-turn attacks, plays a crucial role in uncovering underlying security vulnerabilities. However, prior methods, including sophisticated multi-turn approaches, often struggle to adapt to the evolving dynamics of dialogue as interactions progress. To address this challenge, we propose GRAF(JailBreaking via Globally Refining and Adaptively Fabricating), a novel multi-turn jailbreaking method that globally refines the attack trajectory at each interaction. In addition, we actively fabricate model responses to suppress safety-related warnings, thereby increasing the likelihood of eliciting harmful outputs in subsequent queries. Extensive experiments across six state-of-the-art LLMs demonstrate the superior effectiveness of our approach compared to existing single-turn and multi-turn jailbreaking methods. Our code will be released at https://github.com/Ytang520/Multi-Turn_jailbreaking_Global-Refinment_and_Active-Fabrication.
@@ 1 Introduction
Large Language Models (LLMs) have recently exhibited superior performance across a wide range of tasks, including role-playing, information retrieval, and mathematical reasoning (Kim et al., 2024; Cheong et al., 2024; Tang et al., 2023; Guo et al., 2025; Hurst et al., 2024). Despite these numerous beneficial applications, the powerful capabilities of LLMs can also be exploited for malicious purposes (Gupta et al., 2023; Chao et al., 2023; Yi et al., 2024a). As a result, a growing body of research has focused on developing safety-enhancing mechanisms aimed at preventing models from following harmful instructions (Xu et al., 2024; Inan et al., 2023; Zeng et al., 2024a).
Figure 1: (a): GRAF achieves the highest attack success rate compared with baselines on Harmbench (Mazeika et al., 2024) evaluated through GPT-Judge (Qi et al., 2023). (b): An example of multi-turn jailbreaking with ChatGPT, compared with single-turn jailbreaking.
Among safety-enhancing efforts, jailbreaking serves as a critical red-team technique that aims to bypass safeguards and expose vulnerabilities in LLMs, thereby informing defenses and improving model security (Liu et al., 2023; Chao et al., 2023; Li et al., 2024b). Recent jailbreaking methods can be broadly classified into two categories. (1) Single-turn jailbreaking (Liu et al., 2023; Chao et al., 2023; Li et al., 2024b; Yi et al., 2024a) attempts to elicit a harmful response in a single interaction. For example, Zou et al. (2023) optimizes suffix texts to increase the probability of harmful outputs, and Ding et al. (2023) hides malicious intents by rephrasing prompts and embedding them within benign task contexts. These single-turn approaches often rely on explicit malicious content and thus can be detected and defended against by target models; alternative single-turn strategies try to obscure malicious intent through rephrasing or complex logical chains, but such obfuscation may instead produce irrelevant or off-target replies (Souly et al., 2024; Huang et al., 2024; Russinovich et al., 2024). (2) Multi-turn jailbreaking introduces an auxiliary attacker model that conducts a sequence of dialogue turns to gradually circumvent safeguards (Russinovich et al., 2024; Li et al., 2024a; Gibbs et al., 2024; Ren et al., 2024). Concretely, the attacker model constructs a series of queries $Q=\{q_{1},q_{2},\dots,q_{N}\}$ (the jailbreaking trajectory) that conceals the malicious intent while progressively steering the target model toward a successful jailbreak (see Figure 1). However, most existing multi-turn jailbreaking methods rely on pre-defined templates to construct the jailbreaking trajectory  (Yang et al., 2024a; Jiang et al., 2024b), keeping the queries fixed throughout the entire attack (Cheng et al., 2024). Other approaches attempt to refine queries during the attack, but these refinements are usually limited to local updates of the next query. Such incremental adjustments may lead to off-topic interactions and are often insufficient to achieve a successful jailbreak within the restricted number of attempts (Ren et al., 2024; Russinovich et al., 2024). Furthermore, prior methods typically accept the target model’s intermediate responses without modification and simply append them to the dialogue history. This practice may propagate irrelevant or unhelpful content, thereby reducing the effectiveness of subsequent queries and ultimately resulting in suboptimal performance (Yang et al., 2024b).
@@
To this end, we propose a novel multi-turn jailbreaking method, named GRAF(JailBreaking via Globally Refining and Adaptively Fabricating), which globally refines the jailbreaking trajectory and actively fabricates the dialogue history throughout the multi-turn process. We begin by asking the attacker model to initialize a comprehensive jailbreaking trajectory for the given task. At each turn $i$, the query $q_{i}$ from the trajectory $Q$ is sent to the target model to obtain the corresponding answer $a_{i}$. If the target model refuses to answer $q_{i}$, the attacker model revises $q_{i}$ and retries until a non-rejective response is obtained or the maximum number of retries is reached. Once a non-rejective answer is returned, we globally refine all remaining queries $Q_{>i}$ based on the dialogue history up to $a_{i}$. In contrast to prior work that performs only local updates, refining all subsequent queries at each step helps the attack remain focused on the ultimate jailbreak goal and avoid producing off-topic or ineffective queries. We further introduce an active fabrication mechanism with two key strategies: (1) if the target model persistently rejects $q_{i}$, we discard the pair $(q_{i},a_{i})$ and move on to $q_{i+1}$; (2) when a non-rejective answer $a_{i}$ is returned, we proactively modify it by removing safety-related warnings before appending it to the dialogue history. These strategies improve the attacker model’s ability to recover from rejections and to maintain effective progress toward a successful jailbreak. The process repeats until the final query $q_{N}$ receives a valid response.
@@
  • We introduce a novel multi-turn jailbreaking method that globally refines the jailbreaking trajectory through updating subsequent queries inside the jailbreaking trajectory to adapt to the context, and actively fabricates the dialogue history to enhance the jailbreaking success.
  • We perform extensive analyses of multi-turn jailbreaking methods. We find that our approach shifts the representations of harmful queries toward those of harmless queries, which helps explain the effectiveness of multi-turn attacks.
@@ 2.1 Responsible Large Language Models
Large Language Models have demonstrated remarkable capabilities across diverse domains (Wei et al., 2022a; b; Kim et al., 2024; Cheong et al., 2024; Tang et al., 2023). However, due to inappropriate or unfiltered data in training corpora, they can produce responses that violate human values (Cao et al., 2023; Zou et al., 2023; Yi et al., 2024a). Accordingly, significant effort has been devoted to aligning LLMs with human values (Inan et al., 2023; Rafailov et al., 2023; Korbak et al., 2023; Ji et al., 2023). Within these efforts, jailbreaking—designed to probe LLMs’ safety mechanisms—plays a crucial role in exposing underlying vulnerabilities and informing improvements to model safety (Zou et al., 2023; Yi et al., 2024a; Chao et al., 2023).
@@ 2.2 Jailbreaking
Mainstream attacks primarily focus on single-turn jailbreaks, which allow only a single attack on the target model during the test phase (Zou et al., 2023; Yao et al., 2024; Yi et al., 2024a), like using optimization-based methods to generate adversarial prompts (Zou et al., 2023; Liu et al., 2023) or manipulate the prompt itself to achieve jailbreaks (Ding et al., 2023; Zeng et al., 2024b; Chao et al., 2023). Most of these methods include explicit malicious content to convey their intent to the target model, and can be easily defended by those input filters (Russinovich et al., 2024). Some methods attempt to obscure these phrases by rephrasing the harmful queries (Huang et al., 2024; Souly et al., 2024); these attempts can confuse the LLM, leading it to generate general or even irrelevant responses that fail to fulfill the user’s intent.
@@
Given the real-world scenario of multiple interactions with LLMs, some researchers have extended attacks to multi-turn jailbreaking. Multi-turn jailbreaking allows multiple attacks on the target model, with its final response during the dialogue evaluated for the success or failure of the attack (Russinovich et al., 2024; Gibbs et al., 2024; Li et al., 2024a; Ren et al., 2024; Yang et al., 2024a; Jiang et al., 2024b). Therefore, malicious intent can be concealed within the jailbreaking trajectory (Russinovich et al., 2024; Ren et al., 2024), rather than being explicitly stated in the jailbreaking queries. For instance, Russinovich et al. (2024) carefully designs a series of queries to avoid explicitly expressing harmful content in the final query, while Wang et al. (2024); Sun et al. (2024) iteratively refines queries to gradually align query semantics with user intent. However, these approaches often rely on fixed templates (Yang et al., 2024a; Jiang et al., 2024b). Some methods (Ren et al., 2024; Russinovich et al., 2024) attempt to refine queries during the attack, but they typically make only local updates by modifying the immediate next query, which can result in off-topic queries and may fail to jailbreak. In addition, previous approaches often directly append the target model’s responses to the dialogue history without any revision, which may hinder the model from answering subsequent queries and lead to suboptimal performance.
@@
Our approach differs from prior work in two main ways. First, we globally refine the jailbreaking trajectories by updating all the subsequent queries at once to adapt to the dynamic dialogue history. In contrast,  Ren et al. (2024) only refines the query when it is rejected by the target model, and  Russinovich et al. (2024) generates the next query based solely on the previous turn. Second, instead of simply appending the target model’s responses to the dialogue history, we detect and remove safety warnings, which helps the attacker model to smoothly progress toward jailbreaking success.
3 Methodology
@@
In this section, we present our proposed approach, GRAF. We first describe the initialization of the jailbreaking trajectory in Subsection 3.1. We then introduce the Global Refinement in Subsection 3.2, followed by the Active Fabrication techniques in Subsection 3.3. The overall workflow of our method is illustrated in Figure 2.
@@
Figure 2: The workflow of GRAF. The attacker model first initialized a jailbreaking trajectory based on the jailbreaking task (denoted by the Gray arrow). For Global Refinement, at each turn, the query $q_{i}$ in the jailbreaking trajectory $Q$ is first fed into the target model to obtain the answer $a_{i}$. If the query $q_{i}$ is refused, the attacker model iteratively revises $q_{i}$. If a successful response is not obtained after the maximum number of retries, the model then revises all subsequent queries $Q_{\geq i}$ to redirect the current trajectory. Once a non-rejected response is returned, the attacker model globally refines all following queries $Q_{>i}$ using the dialogue history up to $a_{i}$ (denoted by the Red Arrow). For Active Fabrication, if $q_{i}$ continues to be rejected, the $(q_{i},a_{i})$ pair is discarded, and the model proceeds to $q_{i+1}$; In addition, any non-rejected response $a_{i}$ is post-processed to remove safety warnings (denoted by the Blue Arrow).
3.1 Trajectory Initialization
Since LLMs may refuse to generate a jailbreaking trajectory when directly prompted, we follow Ren et al. (2024) and first query the attacker model for task-relevant information. For instance, if the malicious task is to obtain instructions for constructing a bomb, the attacker model might return information about a known criminal, such as Ramzi Yousef, who has historically built bombs for terrorism. Based on this information, the attacker model begins with an innocuous initial query (e.g., What has Ramzi Yousef done in history?) and then issues a sequence of queries that gradually guide the target model’s responses toward the jailbreaking goal. Implementation details are provided in Appendix A.1, and example initialized trajectories are shown in Appendix G.6.1.
@@ 3.2 Global Refinement
As shown in Figure 2, during each turn $i$, the query $q_{i}$ in the jailbreaking trajectory $Q=\{q_{1},q_{2},...,q_{n}\}$ is first fed into the target model to obtain the corresponding answer $a_{i}$. Despite the efforts to conceal malicious intent within the jailbreaking trajectory, some queries may still contain harmful content and be rejected by the target model. If the query $q_{i}$ is rejected, we first revise it by replacing the harmful content with referential expressions such as ’that in your previous response’. This helps to conceal explicit malicious content within the dialogue history while preserving the core intent of each query. However, the target model may still reject the query $q_{i}$ after the pre-defined maximum number of tries. We then revise all subsequent queries $Q_{\geq i}$ to redirect the current jailbreaking trajectory, which changes the core intent of $q_{i}$ and further conceals its malicious intent.
Once the non-rejected answer $a_{i}$ is obtained, the attacker model globally refines the remaining queries $Q_{>i}=\{q_{i+1},...,q_{N}\}$ based on the dialogue history $\{q_{1},a_{1},...,q_{i},a_{i}\}$ up to $a_{i}$. The refinement helps discover more effective and stealthy jailbreaking trajectories given the current dialogue history. Furthermore, unlike previous approaches (Russinovich et al., 2024; Yang et al., 2024b), our method reduces the likelihood of generating off-topic or ineffective queries. It does so by updating all subsequent queries at each step, keeping the trajectory focused on achieving a successful jailbreak. Therefore, our method increases the probability that the trajectory reaches the target within the allowed number of interactions. To support this process, we monitor the final query in the generated trajectory to check whether its topic deviates from the jailbreaking task. If a deviation is detected, we iteratively refine the trajectory until it becomes relevant or the maximum number of attempts is reached.